# Conceptualization of this study, Methodology, Software

### v2

@@ Abstract
Uncovering new therapeutic uses of existing drugs, drug repositioning offers a fast and cost-effective strategy and holds considerable significance in the realm of drug discovery and development. In recent years, deep learning techniques have emerged as powerful tools in drug repositioning due to their ability to analyze large and complex datasets. However, many existing methods focus on extracting feature information from nearby nodes in the network to represent drugs and diseases, without considering the potential inter-relationships between the features of drugs and diseases, which may lead to inaccurate representations. To address this limitation, we use two features (similarity and association) to capture the potential relationships between the features of drugs and diseases, proposing a Dual-Feature Drug Repositioning Neural Network (DFDRNN) model. DFDRNN uses a self-attention mechanism to extract neighbor features and incorporates two dual-feature extraction modules: the intra-domain dual-feature extraction (IntraDDFE) module for extracting features within a single domain (drugs or diseases) and the inter-domain dual-feature extraction (InterDDFE) module for extracting features across domains. By utilizing these modules, we ensure more appropriate encoding of drugs and diseases. Additionally, a cross-dual-domain decoder is designed to predict drug-disease associations in both domains. Our proposed DFDRNN model outperforms six state-of-the-art methods on four benchmark datasets, achieving an average AUROC of 0.946 and an average AUPR of 0.597. Case studies on two diseases show that the proposed DFDRNN model can be applied in real-world scenarios, demonstrating its significant potential in drug repositioning.
keywords:
[…]
1 Introduction
@@
In the field of medicine, drug development is a complex and time-consuming process. It involves three main stages: the discovery stage, the preclinical stage, and the clinical stage. This process typically takes 10-15 years and around \$2.8 billion on average (Fisher Wilson, 2006; Vijayan et al., 2022). Despite continuous efforts and innovation, the traditional drug development model faces significant risks and sometimes leading to the termination of projects due to concerns about drug efficacy, safety, or commercial viability. Notably, a high proportion (80-90%) of these projects fail during clinical trials (Vijayan et al., 2022; DiMasi et al., 2010). Therefore, finding a more efficient and precise method for drug development is critically important and valuable.
@@
Computational drug repositioning is a powerful strategy that aims to discover alternative uses of already approved drugs taking into account their potential side effects (Dudley et al., 2011; Wang et al., 2024). This approach uses computational and bioinformatics tools and is more efficient and cost-effective than traditional drug research and development methods. It allows for rapid screening of numerous candidate drugs and accelerates the development of new therapeutic options. Repurposing existing drugs can significantly reduce development costs compared to developing new drugs from scratch. In response to new health crises, such as pandemics, repurposing known drugs can provide a quicker response compared to developing new treatments. This method can also identify new usage for drugs that may be effective for diseases lacking effective treatments and thereby addressing significant health challenges.
@@
Machine learning is commonly used in computational drug repositioning. Traditional techniques like random forest (RF), support vector machine (SVM), and decision trees are often applied for this purpose (Cai et al., 2023). For instance, Peng et al. (2017) introduced a novel drug-target interaction screening framework called PUDTI, which utilized SVM to identify potential drug repositioning targets. Shi et al. (2019) proposed a method for predicting drug-target interactions using Lasso dimensionality reduction and a random forest classifier. Zhao et al. (2022) and Yang et al. (2023) employed random forests in their final prediction step, while other techniques such as decision trees (Xuan et al., 2019), matrix factorization(Yang et al., 2021; Luo et al., 2018; Sadeghi et al., 2022), and logistic regression (Yu et al., 2020) have also been developed.
@@
In recent years, deep learning technology has become a powerful tool in bioinformatics, as evidenced by (Zhang et al., 2017; Mahdaddi et al., 2021; Bansal et al., 2024). Unlike traditional machine learning methods, deep learning technology can effectively capture high-level hidden representation using deep architectures without manual feature selection and tuning (Yu et al., 2022). Existing deep learning methods for drug repositioning often incorporate graph neural networks like Graph Convolutional Networks (GCN) and Graph Attention Networks (GAT) to for model building (Yu et al., 2021; Cai et al., 2021; Sun et al., 2022; Tang et al., 2023; Zhang et al., 2024). For instance, Yu et al. (2021) performed graph convolution operations on the drug-disease heterogeneous network to encode drugs and diseases, but did not account for the differences in encoding from different network spaces. Cai et al. (2021) distinguished between intra-domain and inter-domain feature extraction on the drug-disease heterogeneous network for encoding drugs and diseases. Zhang et al. (2024) extracted features of drugs and diseases from multiple neighborhood spaces and designed a contrastive learning loss to fuse these features. However, these studies did not consider the calculation of information weights from different neighbors. Authors in (Gu et al., 2022; Meng et al., 2024; Huang et al., 2024) used graph attention networks (GAT) to calculate the weights of neighbor aggregation and extract key neighbor information for encoding. Most deep learning models have so far been focused on extracting valuable features from neighborhoods. However, these models often employ single encoding approach and lack intentional encoding and decoding processes for association prediction. This can lead to the introduction of a large amount of noisy information during model training, resulting in the imprecise encoding of drugs and diseases, ultimately affecting the accuracy of prediction results.
@@
In order to address this challenge, we propose a Dual-Feature Drug Repositioning Neural Network (DFDRNN) model that uses two features, similarity, and association, to accurately represent drugs and diseases. In comparison to existing methods, the proposed model incorporates two dual-feature extraction modules: the IntraDDFE module for extracting features within a single domain (drug or disease) and the InterDDFE module for extracting features across domains. DFDRNN uses a self-attention mechanism (SAM) to dynamically adjust the attention level of each input element to neighbor information, providing fine-grained weight allocation. This allows for efficient aggregation of neighbors and enhances the model’s predictive power. Additionally, a cross-dual-domain decoder is designed to predict drug-disease associations in both domains. The key contributions of this paper are:
@@
(1) To predict potential associations between drugs and diseases, we propose a dual-feature drug repositioning neural network (DFDRNN) model. A self-attention mechanism is used capture complex relationship among adjacent nodes. The encoder of DFDRNN consists of two dual-feature extraction modules: the intra-domain dual-feature extraction (IntraDDFE) module and the inter-domain dual-feature extraction (InterDDFE) module.
@@
(2) The IntraDDFE module and InterDDFE module extract features exploiting information within a single domain and across domains, respectively, and track the dynamic changes of dual-feature to realize precise encoding of drugs and diseases.
@@
(3) A cross-dual-domain decoder is proposed to perform cross-computation between two encoded features. This process yields decoding results for both the drug and disease domains, which help effectively predicting the drug-disease associations.
@@
(4) The performance of the proposed method is compared with that of six state-of-the-art methods on four benchmark datasets. The experimental results show that DFDRNN outperforms other models. In addition, case studies on Alzheimer’s and Parkinson’s diseases validated the reliability of DFDRNN in practical applications.
@@
The remainder of this paper is organized as follows. Section 2 describes the datasets and methods, detailing the implementation process of DFDRNN; Section 3 clarifies the experimental design and analyzes the experimental results, and Section 4 concludes this work with future research directions.
2 Materials and Methods
2.1 Datasets
@@
We conducted experiments on four benchmark datasets. The first dataset, Fdataset (Gottlieb et al., 2011), includes 593 drugs from DrugBank (Knox et al., 2024) and 313 diseases listed in the Online Mendelian Inheritance in Man (OMIM) (Hamosh et al., 2005) database, comprising 1933 drug-disease associations. The second dataset, Cdataset (Luo et al., 2016), consists of 663 drugs registered in DrugBank, 409 diseases from the OMIM database, and 2532 known drug-disease associations. The third dataset, Ldataset (Yu et al., 2021), contains 18,416 drug-disease associations between 269 drugs and 598 diseases sourced from the Comparative Toxicogenomics Database(CTD) (Davis et al., 2023). The fourth dataset, LRSSL (Liang et al., 2017), comprises 763 drugs, 681 diseases, and 3051 validated drug-disease associations. Table 1 summarizes the four datasets’ statistics.
@@
This study determines drug similarity based on chemical structures obtained from DrugBank using standardized SMILES notation (Weininger, 1988). The Tanimoto score of the 2D chemical fingerprints of two drugs is then calculated using the chemistry development kit (Steinbeck et al., 2003). Disease similarity, on the other hand, is based on disease phenotypes and is computed using MimMiner (Van Driel et al., 2006). MimMiner gauges disease similarity by analyzing the similarity of MeSH terms in medical descriptions of diseases from the OMIM database (Lipscomb, 2000).
Table 1: The statistics of the four datasets used in this study
@@
2.2 Model structure
Our research is focused on predicting potential associations between drugs and diseases. To this end, we proposed a dual-feature drug repositioning neural network (DFDRNN) model. The encoder of DFDRNN encompasses two dual-feature extraction modules: the intra-domain dual-feature extraction (IntraDDFE) module and the inter-domain dual-feature extraction (InterDDFE) module. While the IntraDDFE module extracts features among drugs (or diseases), the InterDDFE module extracts features between drugs and diseases and changes the features. A self-attention mechanism (SAM) model is employed when aggregating information to capture the complex relationship among adjacent nodes. The decoder of DFDRNN receives the results from the encoder and implements cross-dual-domain decoding, including both the drug and disease domains. Figure 1 illustrates the diagram of the proposed DFDRNN model.
Figure 1: DFDRNN generates the similarity feature and the association feature for drugs and diseases based on established similarity networks and association networks. The encoding process utilizes a multi-layer connected network architecture, with each layer performing IntraDDFE and InterDDFE based on the self-attention mechanism. A layer attention mechanism is also introduced to compute the final similarity and association features. Finally, the decoder calculates drug-disease association scores by decoding separately for the drug and disease domains.
2.2.1 Three networks
@@
Finally, to encode drugs and diseases simultaneously and distinguish IntraDDFE from InterDDFE, we construct two heterogeneous networks $E_{s}$ and $E_{a}$ based on the above matrices $R,D$, and $A$. $E_{s}$ is used for IntraDDFE, while $E_{a}$ is used for InterDDFE, defined as:
@@
2.2.2 Self-attention mechanism
The self-attention mechanism (SAM) can be described as a process associated with three vectors, i.e., one query vector $Q_{i}$, some key vectors $K_{j}$, and some value vector $V_{j}$, which maps $Q_{i}$ and a group of $K_{j}$-$V_{j}$ pairs to a new output vector. The output of the attention is a weighted sum of $V_{j}$ vectors, where the weight for each $V_{j}$ is computed by taking the dot product of the $Q_{i}$ with the corresponding $K_{j}$ (Vaswani et al., 2017). In our IntraDDFE and InterDDFE modules, we use the multi-head SAM to aggregate neighbor information for drugs and diseases. We use the notation $\mathbf{sam}$ to represent the multi-head SAM mapping that involves two parameters $E$ and $H^{in}$, where $E\in\{E_{s},E_{a}\}$ and $H^{in}\in\mathbb{R}^{(n+m)\times k}$ is an input feature matrix with embedding dimension $k$. The specific implementation of $\mathbf{sam}$ can be described as follows:
@@ 2.2.2 Self-attention mechanism
A fully connected layer is applied after the SAM function to extract further features of drugs and diseases. We use the notation $\mathbf{samf}$ to denote the whole process, i.e., $\mathbf{samf}$ is the combination of $\mathbf{sam}$ with a fully connected (FC) layer. Let $w_{r}\in\mathbb{R}^{k\times k}$ and $w_{d}\in\mathbb{R}^{k\times k}$ be trainable parameter matrices, $b_{r}$ and $b_{d}$ be biases, and $\sigma$ be a LeakyRELU (Maas et al., 2013) activation function. Then,
@@
2.2.3 Encoder
The DFDRNN model is a multi-layer connected network architecture, where each layer involves two types of dual-feature extraction modules: IntraDDFE module based on network $E_{s}$ and InterDDFE module based on network $E_{a}$. First, the initial features of drugs and diseases are generated in two features, including the similarity feature $H^{init}_{s}$ and the association feature $H^{init}_{a}$, described as follows.
@@ 2.2.3 Encoder
Now, we consider the process of extracting features in the $\ell$-th layer for $\ell\geq 1$. In the $\ell$-th layer, we denote by $f^{\ell}$ the IntraDDFE function in this layer, $H^{\ell}_{s}$ and $H^{\ell}_{a}$ are the input similarity feature and the input association feature, respectively, $\hat{H}^{\ell+1}_{s}$ and $\hat{H}^{\ell+1}_{a}$ represent the obtained similarity feature and association feature by IntraDDFE, and $\mathbf{samf}^{\ell}_{f}$ the aggregation function of $\ell$-th layer for IntraDDFE. Then,
@@
Since IntraDDFE extracts features within a single domain, the dual-feature does not undergo dynamic changes. Thus, the input similarity features $H^{\ell}_{s}$ produce the output similarity features $\hat{H}^{\ell+1}_{s}$, and the input association features $H^{\ell}_{a}$ produce the output association features $\hat{H}^{\ell+1}_{a}$. Next, we use $g^{\ell}$ to denote the InterDDFE function in the $\ell$-th layer. Similarly, $\tilde{H}^{\ell+1}_{s}$ and $\tilde{H}^{\ell+1}_{a}$ represent the obtained similarity feature and association feature by InterDDFE, respectively, and $\mathbf{samf}^{\ell}_{g}$ represents the aggregation function. Then,
@@
Since the InterDDFE module extracts features across domains, dual-feature undergo dynamic changes. Specifically, the similarity feature $H^{\ell}_{s}$ is changed into the association feature $\tilde{H}^{\ell+1}_{a}$, and the association feature $H^{\ell}_{a}$ is changed into the similarity feature $\tilde{H}^{\ell+1}_{s}$.
@@
As the depth of network layers increases, the valuable information for predicting drug-disease association extracted by the DFDRNN model decreases. Therefore, we introduce a layer attention mechanism to allocate different weights to the results of different layers. We use $L$ to represent the total number of layers in DFDRNN. The weight of the $\ell$-th layer, denoted by $\beta^{\ell}$, is initialized by $1/L$ and automatically learned by the model. Then, the final encoding of the similarity feature and the association feature $H^{r}_{s}$, $H^{r}_{a}$, $H^{d}_{s}$, and $H^{d}_{a}$ of drugs and diseases are obtained by implementing the layer attention mechanism as follows.
@@
2.2.4 Decoder
A cross-dual-domain decoder is utilized to predict drug-disease association. We perform cross computation between two features, using $s_{r}\in\mathbb{R}^{n\times m}$ and $s_{d}\in\mathbb{R}^{n\times m}$ to represent the prediction values for the drug domain and the disease domain, respectively. Let $s\in\mathbb{R}^{n\times m}$ denote the final score matrix of drug-disease associations, where $s_{ij}$ represents the prediction score between drug $r_{i}$ and disease $d_{j}$. That is,

### v3

# Boosting drug-disease association prediction for drug repositioning via dual-feature extraction and cross-dual-domain decoding

@@ Abstract
The extraction of biomedical data has significant academic and practical value in contemporary biomedical sciences. In recent years, drug repositioning, a cost-effective strategy for drug development by discovering new indications for approved drugs, has gained increasing attention. However, many existing drug repositioning methods focus on mining information from adjacent nodes in biomedical networks without considering the potential inter-relationships between the feature spaces of drugs and diseases. This can lead to inaccurate encoding, resulting in biased mined drug-disease association information. To address this limitation, we propose a new model called Dual-Feature Drug Repurposing Neural Network (DFDRNN). DFDRNN allows the mining of two features (similarity and association) from the drug-disease biomedical networks to encode drugs and diseases. A self-attention mechanism is utilized to extract neighbor feature information. It incorporates two dual-feature extraction modules: the single-domain dual-feature extraction (SDDFE) module for extracting features within a single domain (drugs or diseases) and the cross-domain dual-feature extraction (CDDFE) module for extracting features across domains. By utilizing these modules, we ensure more appropriate encoding of drugs and diseases. A cross-dual-domain decoder is also designed to predict drug-disease associations in both domains. Our proposed DFDRNN model outperforms six state-of-the-art methods on four benchmark datasets, achieving an average AUROC of 0.946 and an average AUPR of 0.597. Case studies on two diseases show that the proposed DFDRNN model can be applied in real-world scenarios, demonstrating its significant potential in drug repositioning.
Index Terms:
Drug Repositioning, Self-Attention Mechanism, Dual-Feature Extraction, Cross-Dual-Domain
I Introduction
@@
Drug development is a significant yet complex and time-consuming process. It consists of three main stages: the discovery stage, the preclinical stage, and the clinical stage. Typically, the entire process takes about 10 to 15 years and costs around \$2.8 billion on average [1, 2]. Despite ongoing efforts and innovations, the traditional drug development model faces considerable risks, often resulting in project terminations due to concerns about drug efficacy, safety, or commercial viability. Notably, a high percentage (80-90%) of these projects fail during clinical trials. Consequently, finding a more efficient and precise method for drug development is critically important and valuable.
@@
Given the high cost of drug development, computational drug repositioning has become a powerful strategy for discovering alternative uses of already approved drugs while considering their potential effects [3]. This approach utilizes computational and bioinformatics tools, providing greater efficiency and cost-effectiveness compared to traditional drug research and development methods. It enables the rapid screening of numerous candidate drugs and accelerates the development of new therapeutic options. Repurposing existing drugs can significantly lower development costs when compared to creating new drugs from scratch. In response to health crises, such as pandemics, repurposing known drugs offers a quicker solution than developing entirely new treatments. Furthermore, this method can unveil new applications for drugs that may be effective against diseases that currently lack effective therapies, thereby addressing major health challenges.
@@
Machine learning has been commonly used in computational drug repositioning by information extraction from biomedical data. Techniques like random forest (RF), support vector machine (SVM), and decision trees are often applied for this purpose [4]. For instance, [5] introduced a novel drug-target interaction screening framework called PUDTI, which utilized SVM to identify potential drug repositioning targets. [6] proposed a method for predicting drug-target interactions using Lasso dimensionality reduction and a random forest classifier. [7] employed random forests in their final prediction step, while other techniques such as decision trees [8], matrix factorization[9, 10], and logistic regression [11] have also been developed.
@@
In recent years, deep learning technology has become a powerful tool in bioinformatics, as evidenced by [12, 13, 14, 15]. Unlike traditional machine learning methods, deep learning techniques leverage deep architectures to effectively capture high-level latent representations without requiring manual feature selection and tuning [16]. Existing deep learning methods for drug repositioning often incorporate graph neural networks like Graph Convolutional Networks (GCN) and Graph Attention Networks (GAT) to for model building [17, 18, 19, 20, 21]. For instance, [17] performed graph convolution operations on the drug-disease heterogeneous network to encode drugs and diseases but did not account for the differences in encoding from different network spaces. [18] distinguished intra-domain and inter-domain feature extraction on the drug-disease heterogeneous network for encoding drugs and diseases. [21] extracted features of drugs and diseases from multiple neighborhood spaces and designed a contrastive learning loss to fuse these features. However, these studies did not consider calculating information weights from different neighbors. Authors in [22, 23] used graph attention networks (GAT) to calculate the weights of neighbor aggregation and extract key neighbor information for encoding. Most deep learning methods have primarily focused on extracting valuable features from neighborhoods. However, these approaches often rely on a single encoding method, which does not intentionally account for relation prediction. As a result, a significant amount of noise can arise during model training, hindering the effective extraction of drug-disease association information and ultimately impacting the accuracy of the prediction results.
@@
To address this challenge, we propose a new model called Dual-Feature Drug Repositioning Neural Network (DFDRNN). DFDRNN simultaneously uses two features, similarity and association, to encode drugs and diseases accurately. Compared to existing methods, the proposed model incorporates two dual-feature extraction modules: the SDDFE module for extracting features within a single domain (drug or disease) and the CDDFE module for extracting features across domains. DFDRNN uses a self-attention mechanism (SAM) to dynamically adjust the attention level of each input element to neighbor information, providing fine-grained weight allocation. This allows for efficient aggregation of neighbors and enhances the model’s predictive power. A cross-dual-domain decoder is also designed to predict drug-disease associations in both domains. The key contributions of this work are:
@@
(1) To predict potential associations between drugs and diseases, we propose a dual-feature drug repositioning neural network (DFDRNN) model. A self-attention mechanism is used to capture complex relationships among adjacent nodes. The encoder of DFDRNN consists of two dual-feature extraction modules: the single-domain dual-feature extraction (SDDFE) module and the cross-domain dual-feature extraction (CDDFE) module.
@@
(2) The SDDFE module and the CDDFE module extract information from similarity networks and association networks, respectively, and track the dynamic changes of dual-feature to achieve precise encoding of drugs and diseases.
@@
(3) A cross-dual-domain decoder is proposed to perform cross-computation between two encoded features. This process yields decoding results for both the drug and disease domains, which help effectively predict the drug-disease associations.
@@
(4) The performance of the proposed method is compared with that of six state-of-the-art methods on four benchmark datasets. The experimental results show that DFDRNN outperforms other methods. In addition, case studies on Alzheimer’s and Parkinson’s diseases validated the reliability of DFDRNN in practical applications.
@@
The rest of this paper is structured as follows: Section II introduces the datasets and methods, providing details on the DFDRNN implementation; Section III explains the experimental setup and discusses the results; Section IV concludes the study and suggests future research directions.
II Materials and Methods
II-A Datasets
@@
We performed experiments on four benchmark datasets. The first, Gdataset [24], comprises 593 drugs from DrugBank [25] and 313 diseases from the Online Mendelian Inheritance in Man (OMIM) [26], with a total of 1933 drug-disease associations. The second dataset, Cdataset [27], consists of 663 drugs listed in DrugBank, 409 diseases from OMIM, and 2532 known associations. Ldataset [17], the third dataset, contains 18,416 associations between 269 drugs and 598 diseases, derived from the Comparative Toxicogenomics Database (CTD) [28]. Lastly, LRSSL [29] contains 763 drugs, 681 diseases, and 3051 validated drug-disease associations. A summary of the statistics for these datasets can be found in Table I.
@@
This study determines drug similarity based on chemical structures obtained from DrugBank using standardized SMILES notation [30]. For disease similarity, it is determined based on disease phenotypes and computed via MimMiner [31], which measures similarity by analyzing the MeSH terms in the medical descriptions of diseases found in the OMIM database [32].
TABLE I: Summary statistics of the four datasets
@@
II-B Model structure
Our research is focused on predicting potential associations between drugs and diseases. To this end, we proposed a dual-feature drug repositioning neural network (DFDRNN) model. The encoder of DFDRNN encompasses two dual-feature extraction modules: the single domain dual-feature extraction (SDDFE) module and the cross domain dual-feature extraction (CDDFE) module. While the SDDFE module extracts features among drugs (or diseases), the CDDFE module extracts features between drugs and diseases and changes the features. A self-attention mechanism (SAM) model is employed when aggregating information to capture the complex relationship among adjacent nodes. The decoder of DFDRNN receives the results from the encoder and implements cross-dual-domain decoding, including both the drug and disease domains. Figure 1 illustrates the diagram of the proposed DFDRNN model.
Figure 1: DFDRNN generates the similarity feature and the association feature for drugs and diseases. The encoding process employs a multi-layer connected network architecture, where each layer performs SDDFE and CDDFE, while tracking the dynamic changes of dual-feature. A layer attention mechanism computes final features. The decoder then calculates drug-disease association scores separately for each domain.
II-B1 Three key networks
@@
Finally, to encode drugs and diseases simultaneously and distinguish SDDFE from CDDFE, we construct two heterogeneous networks $E_{s}$ and $E_{a}$ based on the above matrices $R,D$, and $A$. $E_{s}$ is used for SDDFE, while $E_{a}$ is used for CDDFE, defined as:
@@
II-B2 Self-attention mechanism
The self-attention mechanism (SAM) can be described as a process associated with three vectors, i.e., one query vector $Q_{i}$, some key vectors $K_{j}$, and some value vector $V_{j}$, which maps $Q_{i}$ and a group of $K_{j}$-$V_{j}$ pairs to a new output vector. The output of the SAM is a weighted sum of $V_{j}$ vectors, where the weight for each $V_{j}$ is computed by taking the dot product of the $Q_{i}$ with the corresponding $K_{j}$ [33]. In our SDDFE and CDDFE modules, we use the multi-head SAM to aggregate neighbor information for drugs and diseases. We use the notation $\mathbf{sam}$ to represent the multi-head SAM mapping that involves two parameters $E$ and $H^{in}$, where $E\in\{E_{s},E_{a}\}$ and $H^{in}\in\mathbb{R}^{(n+m)\times k}$ is an input feature matrix with embedding dimension $k$. The specific implementation of $\mathbf{sam}$ can be described as follows:
@@ II-B2 Self-attention mechanism
After the SAM function, a fully connected layer is employed to further extract the features of drugs and diseases. We use the notation $\mathbf{samf}$ to denote the whole process, i.e., $\mathbf{samf}$ is the combination of $\mathbf{sam}$ with a fully connected (FC) layer. Let $w_{r}\in\mathbb{R}^{k\times k}$ and $w_{d}\in\mathbb{R}^{k\times k}$ be trainable parameter matrices, $b_{r}$ and $b_{d}$ be biases, and $\sigma$ be a LeakyRELU [34] activation function. Then,
@@
II-B3 Encoder
The DFDRNN model is a multi-layer connected network architecture, where each layer involves two types of dual-feature extraction modules: SDDFE module based on network $E_{s}$ and CDDFE module based on network $E_{a}$. First, the initial features of drugs and diseases are generated in two features, including the similarity feature $H^{init}_{s}$ and the association feature $H^{init}_{a}$, described as follows.
@@ II-B3 Encoder
Now, we consider the process of extracting features in the $\ell$-th layer for $\ell\geq 1$. In the $\ell$-th layer, we denote by $f^{\ell}$ the SDDFE function in this layer, $H^{\ell}_{s}$ and $H^{\ell}_{a}$ are the input similarity feature and the input association feature, respectively, $\hat{H}^{\ell+1}_{s}$ and $\hat{H}^{\ell+1}_{a}$ represent the obtained similarity feature and association feature by SDDFE, and $\mathbf{samf}^{\ell}_{f}$ the aggregation function of $\ell$-th layer for SDDFE. Then,
@@
Since SDDFE extracts features within a single domain, the dual-feature does not undergo dynamic changes. Thus, the input similarity features $H^{\ell}_{s}$ produce the output similarity features $\hat{H}^{\ell+1}_{s}$, and the input association features $H^{\ell}_{a}$ produce the output association features $\hat{H}^{\ell+1}_{a}$. Next, we use $g^{\ell}$ to denote the CDDFE function in the $\ell$-th layer. Similarly, $\tilde{H}^{\ell+1}_{s}$ and $\tilde{H}^{\ell+1}_{a}$ represent the obtained similarity feature and association feature by CDDFE, respectively, and $\mathbf{samf}^{\ell}_{g}$ represents the aggregation function. Then,
@@
Since the CDDFE module extracts features across domains, dual-feature undergo dynamic changes. Specifically, the similarity feature $H^{\ell}_{s}$ is changed into the association feature $\tilde{H}^{\ell+1}_{a}$, and the association feature $H^{\ell}_{a}$ is changed into the similarity feature $\tilde{H}^{\ell+1}_{s}$.
@@
As the depth of network layers increases, the valuable information for predicting drug-disease association extracted by the DFDRNN model decreases. Consequently, a layer attention mechanism is incorporated to assign distinct weights to the outputs of each layer. We use $L$ to represent the total number of layers in DFDRNN. The weight of the $\ell$-th layer, denoted by $\beta^{\ell}$, is initialized by $1/L$ and automatically learned by the model. Then, the final encoding of the similarity feature and the association feature $H^{r}_{s}$, $H^{r}_{a}$, $H^{d}_{s}$, and $H^{d}_{a}$ of drugs and diseases are obtained by implementing the layer attention mechanism as follows.
@@
II-B4 Decoder
A cross-dual-domain decoder is utilized to predict drug-disease association. We perform cross computation between two features, using $s_{r}\in\mathbb{R}^{n\times m}$ and $s_{d}\in\mathbb{R}^{n\times m}$ to represent the prediction values for the drug domain and the disease domain, respectively. Let $s\in\mathbb{R}^{n\times m}$ denote the final score matrix of drug-disease associations, where $s_{ij}$ represents the prediction score between drug $r_{i}$ and disease $d_{j}$. That is